# Weize Chen1  , Jiarui Yuan1∗, Chen Qian1Cheng Yang2 Zhiyuan Liu1, Maosong Sun1 1

### v1

@@ E.3 Training
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})+\alpha\mathcal{L}_{\text{NLL}}(c_{i}^{w},y_{i}^{w}|x_{i})$
$=-\log\sigma\bigg{(}\beta\log\frac{M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{M_{t}(c_{i}^{w},y_{i}^{w}|x_{i})}-\beta\log\frac{M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{M_{t}(c_{i}^{l},y_{i}^{l}|x_{i})}\bigg{)}-\alpha\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{|c_{i}^{w}|+|y_{i}^{w}|}$ (4)
@@
E.4 Hyper Parameters
We conducted six iterations of training for each task. The hyper parameters we used are shown in Table 6. The $\alpha$ and $\beta$ in iDPO section of the table correspond to the $\alpha$ and $\beta$ terms in Eq. 4.
@@ E.4 Hyper Parameters
Appendix F Prompts used in Experiments
@@
2. On finding the final answer, ensure to conclude your communication with ”<A>{answer} </A>”, where ”answer” is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
6. You must begin your response with ”{name}:”.
@@
1. On finding the final answer, ensure to conclude your communication with ”<A>{answer} </A>”, where ”answer” is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
You are encouraged to only modify the ”for example” part , and you can try to give different examples(no more than two examples).
@@
2. On finding the final answer, ensure to conclude your communication with ”<A>{answer} </A>”, where ”answer” is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
6. You must begin your response with ”{name}:”.

### v2

@@ H.3 Training
$\mathcal{L}_{\text{DPO+NLL}}$ $=\mathcal{L}_{\text{DPO}}(c_{i}^{w},y_{i}^{w},c_{i}^{l},y_{i}^{l}|x_{i})$
$\quad+\alpha\mathcal{L}_{\text{NLL}}(c_{i}^{w},y_{i}^{w}|x_{i})$
$=-\log\sigma\bigg{(}\beta\log\frac{M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{M_{t}(c_{i}^{w},y_{i}^{w}|x_{i})}$
$\quad-\beta\log\frac{M_{\theta}(c_{i}^{l},y_{i}^{l}|x_{i})}{M_{t}(c_{i}^{l},y_{i}^{l}|x_{i})}\bigg{)}$
$\quad-\alpha\frac{\log M_{\theta}(c_{i}^{w},y_{i}^{w}|x_{i})}{|c_{i}^{w}|+|y_{i}^{w}|}$
@@
H.4 Hyper Parameters
We conducted six iterations of training for each task. The hyper parameters we used are shown in Table 8. The $\alpha$ and $\beta$ in iDPO section of the table correspond to the $\alpha$ and $\beta$ terms in Section H.3.
@@ H.4 Hyper Parameters
Appendix I Prompts used in Experiments
@@
2. On finding the final answer, ensure to conclude your communication with "<A>{answer} </A>", where "answer" is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
6. You must begin your response with "{name}:".
@@
1. On finding the final answer, ensure to conclude your communication with "<A>{answer} </A>", where "answer" is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
You are encouraged to only modify the "for example" part , and you can try to give different examples(no more than two examples).
@@
2. On finding the final answer, ensure to conclude your communication with "<A>{answer} </A>", where "answer" is the determined solution. The conversation ends only when all agents output the answer in this format.
@@
6. You must begin your response with "{name}:".